OpenStudy (anonymous):

Which of the following lines has a slope of -1/2? x + 2y = 0 x - 2y = 0 -x + 2y = 0

OpenStudy (anonymous):

Slop = -a/b

OpenStudy (anonymous):

so......x + 2y = 0

OpenStudy (anonymous):

oh ok so just turning the 1 into x?

OpenStudy (anonymous):

it is 1x + 2y = 0

OpenStudy (anonymous):

-a/b means the coefficent of x and y